OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 3x + y = 1 y + 4 = 5x What is the resulting equation?

OpenStudy (goformit100):

put the value of x or y of second equation in the first

OpenStudy (anonymous):

y=5x-4 3x=-(5x-4) right

OpenStudy (anonymous):

y=5x-4 substitute 3x+y=1 3x+(5x-4)=1 8x-4=1 8x=5 x=5/8

OpenStudy (anonymous):

so whats the answer